Thomas’ Calculus Early Transcendentals 13ed

Study guide for quiz 08

Quiz problems include both the lecture contents and homework problems.

1. Section 6.3: Study the formula of arclength for curves of the form y = f(x) and x = g(y).

2. Section 6.4: Study the formula of area of surfaces of revolution around the a− axis with the generating curves of the form r = f(a) and a = g(r).

3. Skip the exponential change part. Study the meaning of ’separable differential equations’ and the procedure of solving them.